Calculate the current in a series circuit if the voltage is 12.0 V and there are two resistors in series, one with a 2.00 ohm resistance, and one with a 4.00 ohm resistance.



1
Expert's answer
2021-01-07T11:00:47-0500

The total resistance of two resistors in series is sum of their individual resistances:


"R = R1+R2"

where "R1 = 2\Omega", and "R2 = 4\Omega". Thus:


"R = 2\Omega+4\Omega = 6\Omega"

According to the Ohm's law, the current in a circuit with voltage "U = 12V" and resistance "R = 6\Omega" is:


"I = \dfrac{U}{R} = \dfrac{12V}{6\Omega} = 2A"

Answer. 2A.